Good morning sir,please I want you to help me solve my question. If 5^-x=1\n,find 5^2x.
Thank you.

Answer by Theo(13342)

5^-x = 1/n

since 5^-x is the same as 1/5^x, you get:

1/5^x = 1/n

multiply both sides of this equation by 5^x and multiply both sides of this equation by n to get:

n = 5^x which is the same as 5^x = n.

square both sides of this equation to get:

(5^x)^2 = n^2

since (5^x)^2 = 5^(2x), you get:

5^(2x) = n^2

that's your solution.